y(x,t)=8sin(300t-0.15) where x in m ,y in cm and t in second what is the wavelength,frequency,wavespeed
1
Expert's answer
2013-03-25T09:53:24-0400
We have: y(x,t)=8sin(300t-0.15x)
Because the wave equation is y(x,t)=Asin(wt-kx)
so, frequency=(w/(2*pi))=150/pi
wavelength= 2*pi/k =pi*(13,(3)) (m)
wavespeed= wavelength*frequency=w/k= 2000 (m/s)